When an ion-selective electrode for X was immersed in 0.0520 M XCI, the measured potential was 0.0380 V. What is the concentration of X when the potential is 0.0530 V? Assume that the electrode follows the Nemst equation, the temperature is at 25 C, and that the activity coefficient of X* is 1 Number 0.3M
Answer #1

Given that,

EHCL = 0..0380 V

EX = 0.0530 V

[X+] = 0.0520 M

Let's consider the Nernst equation ,

E = E0 - (0.0591/n)*log Q

E = E0 - (0.0591/n)*log(1/[X+]) . --------(1)

0.0380 = E0 - (0.0591/1)*log(1/0.0520)

0.0380 = E0 - 0.0756

E0 = 0.11

Now, substitute the above value in eq(1)

E = E0 - (0.0591/n)*log(1/[X+])

0.0530 =  0.11 - (0.0591/1)*log(1/[X+])

0.964 = log(1/[X+])

1/[X+] = 100.964

1/[X+] = 9.20

[X+] = 0.108 M